OpenStudy (anonymous):

Jules kicks a soccer ball off the ground and into the air with an initial velocity of 25 feet per second. Assume the starting height of the ball is 0 feet. Approximately what maximum height does the soccer ball reach? 0.8 ft 1.6 ft 9.8 ft 19.6 ft

OpenStudy (anonymous):

All I really need help with is finding out how to solve it.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@uri

OpenStudy (welshfella):

I'm assuming the the ball goes vertically upwards use one of the equations for constant acceleration initial velocity u = 25 ft/s final velocity v ( at maxm height) = 0 ft/s acceleration a due to gravity = -32 ft/s^2 ( negative because motion is against gravity) height s is to be found

OpenStudy (welshfella):

v^2 = u^2 + 2as 0 = 25^2 - 2*32* s solve for s

OpenStudy (welshfella):